OpenStudy (anonymous):

Please help! A bag has 4 green marbles, 3 red, and 2 yellow. What is the probability that you pick a red marble, do not replace it, then pick a yello marble?

OpenStudy (anonymous):

Choices are: 2/15 1/15 1/9 1/10 1/9 Is NOT correct!

OpenStudy (anonymous):

If you take out a red and don't replace it then you now have 4 green, 2 red, and 2 yellow. Probability is the part over the whole. So it's 2 yellow/ 8 = 1/4. ... now I'm confused. That's not an answer choice.

OpenStudy (anonymous):

Yeah! Now you can see why i need help lol.

OpenStudy (anonymous):

I'm correct about the meaning of probability so I'm not sure.

OpenStudy (anonymous):

Yeah! I have no idea

OpenStudy (agent0smith):

A bag has 4 green marbles, 3 red, and 2 yellow. What is the probability that you pick a red marble, do not replace it, then pick a yellow marble? First find the total: 9 marbles. There's 3 reds. Chance of picking a red is thus 3/9. Since we don't replace it, there's now only 8 remaining marbles. There are 2 yellow, so the chance of picking a yellow is 2/8. To find the probability of picking a red and then a yellow, multiply those two probabilities together:\[\frac{ 3 }{ 9 } \times \frac{ 2 }{ 8 } = \]

OpenStudy (anonymous):

OH!!! That makes sense! xD

OpenStudy (anonymous):

It's the probability of doing the two actions in succession. That's why I was wrong.

OpenStudy (agent0smith):

But that gives 1/12, which also isn't an answer choice... even though I'm pretty sure that's the correct solution.

OpenStudy (agent0smith):

Unless there's an error in the question

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

thanks guys!